12. A committee of 3 members is to be formed consisting of one representative each from labor, management,and the public. If there are 3 possible representatives from labor, 2 from management, and 4 from the public, determine how many different committees can be formed?
1
Expert's answer
2020-05-28T18:29:16-0400

"\text{Number of different committees can be formed}\\
=C_{1}^{3} \times C_{1}^{2}\times C_{1}^{4}=3 \times 2\times 4 =24"